OpenStudy (anonymous):

|9x+6|=|x-9| The solution set is

OpenStudy (zarkon):

solve 9x+6=x-9 and 9x+6=-(x-9)

OpenStudy (anonymous):

The soultion set is two numbers. Using commas to separate answers

OpenStudy (zarkon):

correct...solve the two equations I wrote above.